The probability of getting a sum divisible by three if a pair of fair dice is rolled and the numbers that turn up are noted.

1
Expert's answer
2022-02-27T15:03:15-0500

We obtain a total of 36 outcomes by rolling a pair of fair dice.

in 12 out of the 36 possibilities, the sum is divisible by 3. The 12 possibilities are (1,2) (1,5) (2,1) (2,4) (3,3) (3,6) (4,2) (4,5) (5,1) (5,4) (6,3) (6,6)

Thus the probability of getting a sum divisible by three = 12/36 = 1/3